Barron Chemical uses a thermoplastic polymer to enhance the appearance of certain RV panels. The initial cost of one process was $130,000 with annual costs of $48,000. Revenues are $84,000 in year 1, increasing by $1000 per year. A salvage value of $23,000 was realized when the process was discontinued after 8 years. What rate of return did the company make on the process?
Answer #1

Let the rate of retun be I%, then

PW at I% = -130000 + (84000 - 48000)*(P/A, I%,8) + 1000 * (P/G, I%,8) + 23000 * (P/F, I%,8) = 0

-130000 + 36000*(P/A, I%,8) + 1000 * (P/G, I%,8) + 23000 * (P/F, I%,8) = 0

dividing by 1000

36*(P/A, I%,8) + (P/G, I%,8) + 23 * (P/F, I%,8) = 130

using trail and error method

When I = 10%, value of 36*(P/A, I%,8) + (P/G, I%,8) + 23 * (P/F, I%,8) = 218.8156

When I = 15%, value of 36*(P/A, I%,8) + (P/G, I%,8) + 23 * (P/F, I%,8) = 181.5430

When I = 20%, value of 36*(P/A, I%,8) + (P/G, I%,8) + 23 * (P/F, I%,8) = 153.3699

When I = 25%, value of 36*(P/A, I%,8) + (P/G, I%,8) + 23 * (P/F, I%,8) = 131.6465

When I = 26%, value of 36*(P/A, I%,8) + (P/G, I%,8) + 23 * (P/F, I%,8) = 127.9075

using interpolation

I = 25% + (131.6465 - 130) / (131.6465 - 127.9075) * (26% - 25%)

I = 25% + 0.44035%

I = 25.44% (Approx)